# Superlinear Convergence of an Interior Point Algorithm on Linear Semi-definite Feasibility Problems with Application to Linear Matrix Inequalities

### v3

@@ 4 An Infeasible Interior Point Algorithm on the Homogeneous Feasibility Model
[Corrector Step] Find the solution $(\Delta X^{1}_{c},\Delta Y^{1}_{c})$ of the linear system (111), (112), with $\sigma=1-\overline{\alpha}$ and $\overline{r}=0$. Set
$\begin{array}[]{c}X^{1}_{+}=\overline{X}^{1}+\Delta X^{1}_{c},\ \ Y^{1}_{+}=\overline{Y}^{1}+\Delta Y^{1}_{c},\\ \mu_{+}^{1}=(1-\overline{\alpha})\mu^{1}.\end{array}$
@@
$\begin{array}[]{c}X^{1}_{k+1}=X^{1}_{+},\ \ Y^{1}_{k+1}=Y^{1}_{+},\\ \mu_{k+1}^{1}=\mu_{+}^{1}.\end{array}$
@@
$X^{1}_{k}=\left[\begin{array}[]{cc}X_{k}&0\\ 0&\tau_{k}\end{array}\right],\quad Y^{1}_{k}=\left[\begin{array}[]{cc}Y_{k}&0\\ 0&\kappa_{k}\end{array}\right].$ (127)
Consequently, $\mu_{k}^{1}=\mu_{k}$ and $(X^{1}_{k},Y^{1}_{k})\in\mathcal{N}_{1}(\beta_{1},\mu^{1}_{k})$.
Proof: We show (127) holds by induction on $k\geq 0$. It is clear that (127) holds when $k=0$, by the choice of $X^{1}_{0},Y^{1}_{0}$. Suppose (127) holds for $k=k_{0}\geq 0$. Then, at the $(k_{0}+1)^{\rm{th}}$ iteration of Algorithm 4.2, by comparing the system of equations (93)-(97) and (111), (112), it can be seen easily that
$\Delta X^{1}_{p}=\left[\begin{array}[]{cc}\Delta X_{p}&0\\ 0&\Delta\tau_{p}\end{array}\right],\quad\Delta Y^{1}_{p}=\left[\begin{array}[]{cc}\Delta Y_{p}&0\\ 0&\Delta\kappa_{p}\end{array}\right]$
satisfy (111), (112) when $\sigma=0,\overline{r}=r=\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1})$. Furthermore, the step length $\overline{\alpha}$ in the $(k_{0}+1)^{\rm{th}}$ iteration of both algorithms are the same. These lead to
$\overline{X}^{1}=\left[\begin{array}[]{cc}\overline{X}&0\\ 0&\overline{\tau}\end{array}\right],\quad\overline{Y}^{1}=\left[\begin{array}[]{cc}\overline{Y}&0\\ 0&\overline{\kappa}\end{array}\right].$ (133)
With (133), again comparing the system of equations (93)-(97) and (111), (112), it is also easy to see that
$\Delta X^{1}_{c}=\left[\begin{array}[]{cc}\Delta X_{c}&0\\ 0&\Delta\tau_{c}\end{array}\right],\quad\Delta Y^{1}_{c}=\left[\begin{array}[]{cc}\Delta Y_{c}&0\\ 0&\Delta\kappa_{c}\end{array}\right]$
@@
$\mu_{k}^{1}=\frac{{\rm{Tr}}(X^{1}_{k}Y^{1}_{k})}{n_{1}}=\frac{{\rm{Tr}}(X_{k}Y_{k})+\tau_{k}\kappa_{k}}{n+1}=\mu_{k}.$
Finally, by (127), comparing the definition of the neighborhood $\mathcal{N}(\beta,\mu)$ and the neighborhood $\mathcal{N}_{1}(\beta,\mu^{1})$, and that $\mu_{k}=\mu_{k}^{1}$, we see that since $(X_{k},y_{k},Y_{k},\tau_{k},\kappa_{k})\in\mathcal{N}(\beta_{1},\mu_{k})$ (Remark 4.1), we have $(X_{k}^{1},Y_{k}^{1})\in\mathcal{N}_{1}(\beta_{1},\mu_{k}^{1})$.
@@ 4.1 Superlinear Convergence
We show in this subsection that Algorithm 4.1 applied to the homogeneous feasibility model when the initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ is such that $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ is feasible to the dual SDP (8) with $C=0$ leads to superlinear convergence of iterates generated by the algorithm, besides polynomial iteration complexity. First, we state an additional assumption, Assumption 4.1, which is the assumption of strict complementarity, on the primal-dual SDP pair (4)-(8) with $C=0$ that is needed for this result to hold. Note that strict complementarity assumption is generally considered the minimal requirement for superlinear convergence of interior point algorithms, as investigated for example in [16].
@@
A consequence of the above assumption on the homogeneous feasibility model (16)-(19) is that it has a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ with $X^{\ast}+Y^{\ast}\in S^{n}_{++}$ and $\tau^{\ast}+\kappa^{\ast}>0$. This then implies that its SDLCP representation has a solution $(X^{1,\ast},Y^{1,\ast})$ such that $X^{1,\ast}+Y^{1,\ast}\in S^{n_{1}}_{++}$, that is, the SDLCP representation has a strictly complementary solution.
@@
Given an initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ to Algorithm 4.1, with $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ feasible to dual SDP (8) with $C=0$. Iterates generated by Algorithm 4.1 converge superlinearly in the sense of (135).
Proof: Let us show the result in the theorem by considering iterates generated by Algorithm 4.2 instead. These iterates are related to those generated by Algorithm 4.1 in a close way as shown in Proposition 4.1. We note that Algorithm 4.2 is Algorithm 4.1 in [26], and Assumptions 2.1 and 3.1 in [26] are satisfied for the SDLCP representation of the homogeneous feasibility model (16)-(19) (Proposition 3.2 and strict complementarity). Hence, results in [26] are applicable to our SDLCP representation. The SDLCP representation that Algorithm 4.2 is solving has the structure of an LSDFP (Remark 3.3), and therefore Theorem 5.1888This theorem is reproduced in our context as Theorem A.1 in Appendix A. in [26] can be applied on our SDLCP representation provided that Condition (52) in the theorem ((155) in Appendix A) is satisfied.
Our choice of initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})$ to Algorithm 4.1 leads to an initial iterate, $(X^{1}_{0},Y^{1}_{0})$, to Algorithm 4.2 that satisfies $\mathcal{B}_{1}(Y^{1}_{0})=0$ except possibly the $(m+n+1)^{th}$ entry [^9] Remark 3.2 and Appendix A. of $\mathcal{B}_{1}(Y^{1}_{0})$. Therefore, Condition (52) of Theorem 5.1 in [26] is satisfied, and by the theorem, we have superlinear convergence in the sense that
$\frac{\mu^{1}_{k+1}}{\mu^{1}_{k}}\rightarrow 0,\ \ {\rm{as}}\ k\rightarrow\infty.$
@@
Similar result as Theorem 4.1 also holds when the HKM search direction or the NT search direction is used in Algorithm 4.1 instead of the dual HKM search direction. The equivalent algorithm on the SDLCP representation for the NT search direction is Algorithm 1 in [27]. We can then apply Theorem 4 or Corollary 1 in [27] to conclude superlinear convergence of iterates when the initial iterate to Algorithm 4.1 with the NT search direction is from a strictly feasible solution to the dual SDP (8) with $C=0$. The process to show this is analogous to what we have discussed and we will not repeat it here again.
@@ Appendix A Appendix
(see Theorem 5.1 in [26]) Assume that there exists a solution $(X^{1,\ast},Y^{1,\ast})\in S^{n_{1}}_{+}\times S^{n_{1}}_{+}$ to an LSDFP (expressed as SDLCP (20)-(22)) such that $X^{1,\ast}+Y^{1,\ast}\in S^{n_{1}}_{++}$. Let $(X^{1}_{k},Y^{1}_{k})$ be iterates generated by Algorithm 4.2 on the LSDFP with the initial iterate $(X^{1}_{0},Y^{1}_{0})$ satisfying
$\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&((\mathcal{A}_{1})_{j_{1}+j_{2}+1})_{22}&0\\ 0&0&0\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&((\mathcal{A}_{1})_{m_{1}})_{22}&0\\ 0&0&0\end{array}\right)^{T}\end{array}\right){\rm{svec}}(X_{0}^{1})=0$ (145)
@@
$\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{11}&0&((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{13}\\ 0&0&0\\ ((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{13}^{T}&0&((\mathcal{B}_{1})_{k_{1}+k_{2}+1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{11}&0&((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{13}\\ 0&0&0\\ ((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{13}^{T}&0&((\mathcal{B}_{1})_{\tilde{n}_{1}-m_{1}})_{33}\end{array}\right)^{T}\end{array}\right){\rm{svec}}(Y_{0}^{1})=0,$ (155)
@@
To put things in perspective, under the assumption that there exists a solution $(X^{1,\ast},Y^{1,\ast})\in S^{n_{1}}_{+}\times S^{n_{1}}_{+}$ to the LSDFP (expressed as SDLCP (20)-(22)) such that $X^{1,\ast}+Y^{1,\ast}\in S^{n_{1}}_{++}$, we can assume that $X^{1,\ast}$ and $Y^{1\ast}$ are given by
$X^{1,\ast}=\left(\begin{array}[]{ccc}\Lambda^{\ast}_{11}&0&0\\ 0&0&0\\ 0&0&\lambda_{k_{0}}^{\ast}\end{array}\right),\quad Y^{1,\ast}=\left(\begin{array}[]{ccc}0&0&0\\ 0&\Lambda^{\ast}_{22}&0\\ 0&0&0\end{array}\right),$
where $\Lambda^{\ast}_{11}={\rm{Diag}}(\lambda_{1}^{\ast},\ldots,\lambda^{\ast}_{k_{0}-1})\in S^{k_{0}-1}_{++}$ and $\Lambda^{\ast}_{22}={\rm{Diag}}(\lambda_{k_{0}+1}^{\ast},\ldots,\lambda^{\ast}_{n_{1}})\in S^{n_{1}-k_{0}}_{++}$. Here, $\lambda_{1}^{\ast},\ldots,\lambda_{n_{1}}^{\ast}$ are real numbers greater than zero. In this appendix, we always partition a matrix $S\in S^{n_{1}}$ in the way $X^{1,\ast}$ and $Y^{1,\ast}$ are partitioned, i.e., $S$ is partitioned as $\left(\begin{array}[]{ccc}S_{11}&S_{12}&S_{13}\\ S_{12}^{T}&S_{22}&S_{23}\\ S_{13}^{T}&S_{23}^{T}&S_{33}\end{array}\right)$, where $S_{11}\in S^{k_{0}-1},S_{22}\in S^{n_{1}-k_{0}},S_{33}\in\Re$ and $S_{12}\in\Re^{(k_{0}-1)\times(n_{1}-k_{0})},S_{13}\in\Re^{k_{0}-1},S_{23}\in\Re^{n_{1}-k_{0}}$. Using this partition, we perform block Gaussian eliminations and row operations on $\mathcal{A}_{1},\mathcal{B}_{1}:S^{n_{1}}\rightarrow\Re^{\tilde{n}_{1}}$ in (20) so that they can be written in matrix form as:
$\mathcal{A}_{1}=\left(\begin{array}[]{c}\bar{\mathcal{A}}_{1}\\ 0\end{array}\right),\quad\mathcal{B}_{1}=\left(\begin{array}[]{c}0\\ \bar{\mathcal{B}}_{1}\end{array}\right),$

### v4

@@ 4 An Infeasible Interior Point Algorithm on the Homogeneous Feasibility Model
[Corrector Step] Find the solution $(\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{c},\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{c})$ of the linear system (104), (105), with ${{\color[rgb]{0,0,0}\hat{X}=\overline{\hat{X}},\hat{Y}=\overline{\hat{Y}}}},\sigma=1-\overline{\alpha}$ and $\overline{r}=0$. Set
$\begin{array}[]{c}{{\color[rgb]{0,0,0}\hat{X}}}_{+}={{\color[rgb]{0,0,0}\overline{\hat{X}}}}+\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{c},\ \ {{\color[rgb]{0,0,0}\hat{Y}}}_{+}={{\color[rgb]{0,0,0}\overline{\hat{Y}}}}+\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{c},\\ {{\color[rgb]{0,0,0}\hat{\mu}}}_{+}=(1-\overline{\alpha}){{\color[rgb]{0,0,0}\hat{\mu}}}.\end{array}$
@@
$\begin{array}[]{c}{{\color[rgb]{0,0,0}\hat{X}}}_{k+1}={{\color[rgb]{0,0,0}\hat{X}}}_{+},\ \ {{\color[rgb]{0,0,0}\hat{Y}}}_{k+1}={{\color[rgb]{0,0,0}\hat{Y}}}_{+},\\ {{\color[rgb]{0,0,0}\hat{\mu}}}_{k+1}={{\color[rgb]{0,0,0}\hat{\mu}}}_{+}.\end{array}$
We see that Algorithm 4.2 is similar to Algorithm 4.1. We present the two algorithms in full above as the proof of Proposition 4.1 requires us to make a comparison between them. Having details of both algorithms presented serves the purpose to improve readability of the proof of the proposition.
@@
${{\color[rgb]{0,0,0}\hat{X}}}_{k}=\left[\begin{array}[]{cc}X_{k}&0\\ 0&\tau_{k}\end{array}\right],\quad{{\color[rgb]{0,0,0}\hat{Y}}}_{k}=\left[\begin{array}[]{cc}Y_{k}&0\\ 0&\kappa_{k}\end{array}\right].$ (120)
Consequently, ${{\color[rgb]{0,0,0}\hat{\mu}}}_{k}=\mu_{k}$ and $({{\color[rgb]{0,0,0}\hat{X}}}_{k},{{\color[rgb]{0,0,0}\hat{Y}}}_{k})\in\mathcal{N}_{1}(\beta_{1},{{\color[rgb]{0,0,0}\hat{\mu}}}_{k})$.
Proof: We show (120) holds by induction on $k\geq 0$. It is clear that (120) holds when $k=0$, by the choice of ${{\color[rgb]{0,0,0}\hat{X}}}_{0},{{\color[rgb]{0,0,0}\hat{Y}}}_{0}$. Suppose (120) holds for $k=k_{0}\geq 0$. Then, at the $(k_{0}+1)^{\rm{th}}$ iteration of Algorithm 4.2, by comparing the system of equations (86)-(90) and (104), (105), it can be seen easily that
$\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{p}=\left[\begin{array}[]{cc}\Delta X_{p}&0\\ 0&\Delta\tau_{p}\end{array}\right],\quad\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{p}=\left[\begin{array}[]{cc}\Delta Y_{p}&0\\ 0&\Delta\kappa_{p}\end{array}\right]$
satisfy (104), (105) when $\sigma=0,\overline{r}=r={{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})$. Furthermore, the step length $\overline{\alpha}$ in the $(k_{0}+1)^{\rm{th}}$ iteration of both algorithms are the same. These lead to
${{\color[rgb]{0,0,0}\overline{\hat{X}}}}=\left[\begin{array}[]{cc}\overline{X}&0\\ 0&\overline{\tau}\end{array}\right],\quad{{\color[rgb]{0,0,0}\overline{\hat{Y}}}}=\left[\begin{array}[]{cc}\overline{Y}&0\\ 0&\overline{\kappa}\end{array}\right]$ (126)
in the two algorithms. With (126), again comparing the system of equations (86)-(90) and (104), (105), it is also easy to see that
$\Delta{{\color[rgb]{0,0,0}\hat{X}}}_{c}=\left[\begin{array}[]{cc}\Delta X_{c}&0\\ 0&\Delta\tau_{c}\end{array}\right],\quad\Delta{{\color[rgb]{0,0,0}\hat{Y}}}_{c}=\left[\begin{array}[]{cc}\Delta Y_{c}&0\\ 0&\Delta\kappa_{c}\end{array}\right]$
@@
${{\color[rgb]{0,0,0}\hat{\mu}}}_{k}=\frac{{\rm{Tr}}({{\color[rgb]{0,0,0}\hat{X}}}_{k}{{\color[rgb]{0,0,0}\hat{Y}}}_{k})}{n_{1}}=\frac{{\rm{Tr}}(X_{k}Y_{k})+\tau_{k}\kappa_{k}}{n+1}=\mu_{k}.$
Finally, by (120), comparing the definition of the neighborhood $\mathcal{N}(\beta,\mu)$ and the neighborhood $\mathcal{N}_{1}(\beta,{{\color[rgb]{0,0,0}\hat{\mu}}})$, and that $\mu_{k}={{\color[rgb]{0,0,0}\hat{\mu}}}_{k}$, we see that since $(X_{k},y_{k},Y_{k},\tau_{k},\kappa_{k})\in\mathcal{N}(\beta_{1},\mu_{k})$ (Remark 4.1), we have $({{\color[rgb]{0,0,0}\hat{X}}}_{k},{{\color[rgb]{0,0,0}\hat{Y}}}_{k})\in\mathcal{N}_{1}(\beta_{1},{{\color[rgb]{0,0,0}\hat{\mu}}}_{k})$.
@@ 4.1 Superlinear Convergence
We show in this subsection that Algorithm 4.1 applied to the homogeneous feasibility model of sLSDFP when the initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ is such that $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ is feasible to the sDSDP leads to superlinear convergence of iterates generated by the algorithm, besides polynomial iteration complexity. First, we state an additional assumption, Assumption 4.1, which is the assumption of strict complementarity, on the primal-dual SDP pair (4)-(8) with $C=0$ that is needed for this result to hold. Note that strict complementarity assumption is generally considered the minimal requirement for superlinear convergence of interior point algorithms, as investigated for example in [16].
@@
A consequence of the above assumption on the homogeneous feasibility model (16)-(19) is that it has a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ with $X^{\ast}+Y^{\ast}\in S^{n}_{++}$ and $\tau^{\ast}+\kappa^{\ast}>0$. This then implies that its SDLCP representation, sSDLCP, has a solution $({{\color[rgb]{0,0,0}\hat{X}^{\ast}}},{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}})$ such that ${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}+{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}\in S^{n_{1}}_{++}$, that is, the sSDLCP has a strictly complementary solution.
@@
Given an initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})\in\mathcal{N}(\beta_{1},\mu_{0})$ to Algorithm 4.1, with $(y_{0},Y_{0})\in\Re^{m}\times S^{n}_{++}$ feasible to the sDSDP. Iterates generated by Algorithm 4.1 converge superlinearly in the sense of (128).
Proof: Let us show the result in the theorem by considering iterates generated by Algorithm 4.2 instead. These iterates are related to those generated by Algorithm 4.1 in a close way as shown in Proposition 4.1. We note that Algorithm 4.2 is Algorithm 4.1 in [26], and Assumptions 2.1 and 3.1 in [26] are satisfied for the SDLCP representation, sSDLCP, of the homogeneous feasibility model (16)-(19) (Proposition 3.2 and strict complementarity). Hence, results in [26] are applicable to our SDLCP representation. The SDLCP representation that Algorithm 4.2 is solving has the structure of an LSDFP (Remark 3.3), and therefore Theorem 5.1999This theorem is reproduced in our context as Theorem A.1 in the appendix. in [26] can be applied on our SDLCP representation provided that Condition (52) in the theorem ((148) in the appendix) is satisfied.
Our choice of initial iterate $(X_{0},y_{0},Y_{0},\tau_{0},\kappa_{0})$ to Algorithm 4.1 leads to an initial iterate, $({{\color[rgb]{0,0,0}\hat{X}}}_{0},{{\color[rgb]{0,0,0}\hat{Y}}}_{0})$, to Algorithm 4.2 that satisfies ^${{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}_{0})=0$ except possibly the $(m+n+1)^{th}$ entry [^10] Remark 3.2 and the appendix. of ^${{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}_{0})$. Therefore, Condition (52) of Theorem 5.1 in [26] is satisfied, and by the theorem, we have superlinear convergence in the sense that
$\frac{{{\color[rgb]{0,0,0}\hat{\mu}}}_{k+1}}{{{\color[rgb]{0,0,0}\hat{\mu}}}_{k}}\rightarrow 0,\ \ {\rm{as}}\ k\rightarrow\infty.$
@@
Similar result as Theorem 4.1 also holds when the HKM search direction or the NT search direction is used in Algorithm 4.1 instead of the dual HKM search direction. The equivalent algorithm on the SDLCP representation for the NT search direction is Algorithm 1 in [27]. We can then apply Theorem 4 or Corollary 1 in [27] to conclude superlinear convergence of iterates when the initial iterate to Algorithm 4.1 with the NT search direction is from a strictly feasible solution to the sDSDP. The process to show this is analogous to what we have discussed and we will not repeat it here again.
@@ Appendix A Appendix
(see Theorem 5.1 in [26]) Assume that there exists a solution $({{\color[rgb]{0,0,0}\hat{X}^{\ast}}},{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}})\in S^{n_{1}}_{+}\times S^{n_{1}}_{+}$ to an LSDFP (expressed as SDLCP (20)-(22)) such that ${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}+{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}\in S^{n_{1}}_{++}$. Let $({{\color[rgb]{0,0,0}\hat{X}}}_{k},{{\color[rgb]{0,0,0}\hat{Y}}}_{k})$ be iterates generated by Algorithm 4.2 on the LSDFP with the initial iterate $({{\color[rgb]{0,0,0}\hat{X}}}_{0},{{\color[rgb]{0,0,0}\hat{Y}}}_{0})$ satisfying
$\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{j_{1}+j_{2}+1})_{22}&0\\ 0&0&0\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}0&0&0\\ 0&(({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}})_{m_{1}})_{22}&0\\ 0&0&0\end{array}\right)^{T}\end{array}\right){\rm{svec}}({{\color[rgb]{0,0,0}\hat{X}}}_{0})=0$ (138)
@@
$\left(\begin{array}[]{l}{\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{11}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{13}\\ 0&0&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{13}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{k_{1}+k_{2}+1})_{33}\end{array}\right)^{T}\\ \vdots\\ {\rm{svec}}\left(\begin{array}[]{ccc}(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{11}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{13}\\ 0&0&0\\ (({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{13}^{T}&0&(({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}})_{\tilde{n}_{1}-m_{1}})_{33}\end{array}\right)^{T}\end{array}\right){\rm{svec}}({{\color[rgb]{0,0,0}\hat{Y}}}_{0})=0,$ (148)
@@
To put things in perspective, under the assumption that there exists a solution $({{\color[rgb]{0,0,0}\hat{X}^{\ast}}},{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}})\in S^{n_{1}}_{+}\times S^{n_{1}}_{+}$ to the LSDFP (expressed as SDLCP (20)-(22)) such that ${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}+{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}\in S^{n_{1}}_{++}$, we can assume that ${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}$ and ${{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}$ are given by
${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}=\left(\begin{array}[]{ccc}\Lambda^{\ast}_{11}&0&0\\ 0&0&0\\ 0&0&\lambda_{k_{0}}^{\ast}\end{array}\right),\quad{{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}=\left(\begin{array}[]{ccc}0&0&0\\ 0&\Lambda^{\ast}_{22}&0\\ 0&0&0\end{array}\right),$
where $\Lambda^{\ast}_{11}={\rm{Diag}}(\lambda_{1}^{\ast},\ldots,\lambda^{\ast}_{k_{0}-1})\in S^{k_{0}-1}_{++}$ and $\Lambda^{\ast}_{22}={\rm{Diag}}(\lambda_{k_{0}+1}^{\ast},\ldots,\lambda^{\ast}_{n_{1}})\in S^{n_{1}-k_{0}}_{++}$. Here, $\lambda_{1}^{\ast},\ldots,\lambda_{n_{1}}^{\ast}$ are real numbers greater than zero. In this appendix, we always partition a matrix $S\in S^{n_{1}}$ in the way ${{\color[rgb]{0,0,0}\hat{X}^{\ast}}}$ and ${{\color[rgb]{0,0,0}\hat{Y}^{\ast}}}$ are partitioned, i.e., $S$ is partitioned as $\left(\begin{array}[]{ccc}S_{11}&S_{12}&S_{13}\\ S_{12}^{T}&S_{22}&S_{23}\\ S_{13}^{T}&S_{23}^{T}&S_{33}\end{array}\right)$, where $S_{11}\in S^{k_{0}-1},S_{22}\in S^{n_{1}-k_{0}},S_{33}\in\Re$ and $S_{12}\in\Re^{(k_{0}-1)\times(n_{1}-k_{0})},S_{13}\in\Re^{k_{0}-1},S_{23}\in\Re^{n_{1}-k_{0}}$. Using this partition, we perform block Gaussian eliminations and row operations on ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}},{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}:S^{n_{1}}\rightarrow\Re^{\tilde{n}_{1}}$ in (20) so that they can be written in matrix form as:
${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}=\left(\begin{array}[]{c}{{\color[rgb]{0,0,0}\bar{\mathcal{A}}}}\\ 0\end{array}\right),\quad{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}=\left(\begin{array}[]{c}0\\ {{\color[rgb]{0,0,0}\bar{\mathcal{B}}}}\end{array}\right),$